Find the solution to the system of equations.
x + y = 4 and y = 2x + 1
I NEED HELP PLEASE !

Answer by rfer(16322)

y=2x+1
x+y=4
------------
x+2x+1=4
3x=3
x=1
--------------
1+y=4
y=3
-------------
(1,3)
------------------
y=2x+1
3=2*1+1